Consider the case of Kuhn Co. Kuhn Co. is considering a new project that will require an initial investment of $4 million. It has a target capital structure of 35% debt, 2% preferred stock, and 63% common equity. Kuhn has noncallable bonds outstanding that mature in five years with a face value of $1,000, an annual coupon rate of 10%, and a market price of $1,050.76. The yield on the companys current bonds is a good approximation of the yield on any new bonds that it issues. The company can sell shares of preferred stock that pay an annual dividend of $8 at a price of $92.25 per share. Kuhn does not have any retained earnings available to finance this project, so the firm will have to issue new common stock to help fund it. Its common stock is currently selling for $22.35 per share, and it is expected to pay a dividend of $2.78 at the end of next year. Flotation costs will represent 3% of the funds raised by issuing new common stock. The company is projected to grow at a constant rate of 9.2%, and they face a tax rate of 40%. Determine what Kuhn Companys WACC will be for this project 15.87% 17.46% 14.28% 12.70% Flash Player MAC 32.0.0.114
Answer #1

cost of debt before tax using excel function=RATE(nper,pmt,pv,fv)

=RATE(5,1000*10%,-1050.76,1000)

=8.70%

cost of preferred stock=8/92.25=8.67%

cost of equity=((dividend in next year)/(current price-flotation cost))+growth rate

=(2.78/(22.35-22.35*3%))+9.2%

=22.02%

WACC=63%*22.02%+2%*8.67%+35%*8.70%*(1-40%)

=15.87%

the above is answer..